OpenStudy (anonymous):

find the slope: here are the points (-2,1) (6,-2) so it would be (-2-1) over (6- -2) CORRECT?

OpenStudy (amistre64):

it should be (y1 -y2) over (x1 -x2) (1-(-2)) over (-2 -(6)) = 3/-8 or: -3/8

OpenStudy (anonymous):

and then how woudl i put that into slope intercept equation... y=mx+b

OpenStudy (amistre64):

take either coordinate set, doesnt matter; and plug it into: y = (-3/8)x + b and solve for (b) Lets use (-2,1)... 1 = (-3/8)(-2) + b 1 = (-6/8) + b 1 + (6/8) = b (8/8) + (6/8) = b (14/8) = b we can reduce (14/8) to (7/4) since we know (m) and (b); lets get rid of the values for (x) and (y) to form the equation: y = (-3/8)x + (7/4)

OpenStudy (anonymous):

i just looked at the answer in the book and it said 1/4.... any idea how they got taht?

OpenStudy (amistre64):

yeah; i missed something... (-3/8)(-2) is not a negative, its a positive 1 = (6/8) + b 1 - (6/8) = b (8/8) - (6/8) = b (2/8) = b 2/8 reduces to 1/4 itll get ya everytime :)

OpenStudy (anonymous):

yeah its those simple little things that alwasy get me. i really appreciate all your help!